OpenStudy (anonymous):

Find the center, the radius, the x intercepts and y intercepts for the circle... x^2 + y^2 -8x +6y-11=0

OpenStudy (anonymous):

Complete the square the x and y vaiables independently (x^2-8x+16)+(y^2+6y+9)=11+16+9 Put in standard form (x-4)^2+(y+3)^2=36 Center (4,-3); radius=6 (cont)

OpenStudy (anonymous):

Sorry not working at home. I din't get the line breaks that I expected--I hope you can read it.

OpenStudy (anonymous):

it's ok

OpenStudy (anonymous):

indeed i got that first part but I'm lost after this u.u

OpenStudy (anonymous):

You can get the x-int by letting y=0 and solving for x: (x-4)^2+(0+3)^2=36 or (x-4)^2=27 then square root and add 4 to get: x=4+3sqrt(3) union x=4-3sqrt(3)

OpenStudy (anonymous):

Get y-ints by letting x=0 and solving for y: (0-4)^2+(y+3)^2=36 or (y+3)^2=20; then square root and subtract 3 to get: y=-3+2sqrt(5) union y=-3-2sqrt(5)

OpenStudy (anonymous):

I'm on a school computer and I can't seem to get line breaks, ansd I can't use the equation editor. Sorry!

OpenStudy (anonymous):

its completely fine! dont worry and thank you

OpenStudy (anonymous):

You are welcome